A wave is represented by y = (3.00m) sin Ɵ (2.5t -3.14x). Find the (a) amplitude, (b) angular frequency, (c) frequency, (d) wave number, (e) wavelength, and (f) speed of the wave.



1
Expert's answer
2021-12-20T10:31:03-0500

(a) "A=3.0\ m."

(b) "\omega=2.5\ \dfrac{rad}{s}."

(c)

"\omega=2\pi f,""f=\dfrac{\omega}{2\pi}=\dfrac{2.5\ \dfrac{rad}{s}}{2\pi}=0.4\ Hz."

(d) "k=3.14\ m^{-1}."

(e)

"\lambda=\dfrac{2\pi}{k}=\dfrac{2\pi}{3.14\ m^{-1}}=2\ m."

(f)

"v=f\lambda=0.4\ Hz\times2\ m=0.8\ \dfrac{m}{s}."